OpenStudy (anonymous):

Which of these follows from Newton's laws about the energy of a moving object?

OpenStudy (anonymous):

A. The mechanical energy remains the same if there are only conservative forces acting B. The total of all forms of energy is always constant C. the kinetic energy of an object remains the same if friction is absent D. the total mechanical energy of an object is always constant

OpenStudy (michele_laino):

I think the first option, since we can write the subsequent chain of equalities, for a free falling particle whose mass is m: \[PE = mgh = mg\frac{1}{2}g{t^2} = \frac{1}{2}m{\left( {gt} \right)^2} = \frac{1}{2}m{v^2} = KE\]

OpenStudy (michele_laino):

keep in mind that gravitational force, is a conservative force

OpenStudy (anonymous):

ohh okay!!! thank you:)

OpenStudy (michele_laino):

:)